OpenStudy (anonymous):

Find the circumference of 2 inches.

OpenStudy (kirbykirby):

Is this for a circle with radius 2 inches, or diameter 2 inches?

OpenStudy (kirbykirby):

Then the formula for a circumference of a circle is \[\pi*d=\pi*2=2\pi\] where d = diameter, and that's just pi

OpenStudy (kirbykirby):

or more specifically, \[2\pi\] inches to keep the units

OpenStudy (anonymous):

Just find the area please.

OpenStudy (kirbykirby):

Yes there is: The formula for area is \[2\pi*r^2\] where r = radius Now you need to know that diameter = 2*radius So d = 2r Since d = 2 2 = 2r r = 1 So the area is \[2\pi*r^2=2\pi(1)^2=2\pi \] in inches^2 as units